OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one third x + 7 y = -3x + 3 y = one third x + 3 y = -3x + 7

OpenStudy (anonymous):

OK, so you are given a point and the slope so plug that into y-y1=m(x-x1)

OpenStudy (anonymous):

where m is your slope

OpenStudy (anonymous):

x and y is from the point

OpenStudy (anonymous):

So Im putting it in point slop form?

OpenStudy (anonymous):

yes, to find the equation

OpenStudy (anonymous):

it's currently in standard form

OpenStudy (anonymous):

y = mx+b